agents mulder and scully are driving to georgia to investigate "big blue", a giant aquatic reptile reported to inhabit one of the local lakes. mulder leaves washington at 8:30am and averages 65mph. his partner scully, leaves at 9:00am, following the same path and averaging 68mph. at what time will scully catch up with mulder?
Answer by Paul(988)

Difference between 9:00 and 8:30 = 30 minutes or 30/60 = 0.5h

65(x+0.5)=68x
65x+32.5=68x
3x=32.5
x = 10.83333... round it to 11h.

9:00+11 = 8:00p.m.

Hence he will catch him at 8:00p.m.
Paul.